OpenStudy (kewlgeek555):

A school system is reducing the amount of dumpster loads of trash removed each week. In week 5, there were 60 dumpster loads of waste removed. In week 10, there were 40 dumpster loads removed. Assume that the reduction in the amount of waste each week is linear. Write an equation in function form to show the amount of trash removed each week. a. f(x) = - 4x + 60 b. f(x) = 4x + 60 c. f(x) = - 4x + 80 d. f(x) = 4x + 80

OpenStudy (kewlgeek555):

@agent0smith @ganeshie8 @mathstudent55 @CGGURUMANJUNATH @wio

OpenStudy (anonymous):

i believe it would be C

ganeshie8 (ganeshie8):

you're given two points :- (5, 60) and (10, 40)

ganeshie8 (ganeshie8):

you need to setup the equation in slope-intercept form :- y = mx + b

ganeshie8 (ganeshie8):

start by finding the slope m first

OpenStudy (kewlgeek555):

Well, the slope is positive and the slope is four, correct @ganeshie8 ?

ganeshie8 (ganeshie8):

for this quesiton, YOU saying slope is positive is a crime

ganeshie8 (ganeshie8):

week 5, they dumped 60 week 10, they dumped 40

ganeshie8 (ganeshie8):

clearly, it reduced :o

OpenStudy (kewlgeek555):

Oops...sorry. So, negative?

ganeshie8 (ganeshie8):

(5, 60) and (10, 40) x1 y1 x2 y2 slope = \(\large \frac{y_2-y_1}{x_2-x_1}\)

ganeshie8 (ganeshie8):

slope = \(\large \frac{40-60}{10-5}\) = \(\large \frac{-20}{5}\) = \(\large -4\)

OpenStudy (kewlgeek555):

Okay. So, what would I do next to solve this problem?

ganeshie8 (ganeshie8):

substitute m value in above equation y = mx + b y = -4x + b

ganeshie8 (ganeshie8):

you still need to find b

ganeshie8 (ganeshie8):

use this :- the line goes thru (5, 60)

ganeshie8 (ganeshie8):

so it has to satisfy this point

ganeshie8 (ganeshie8):

y = -4x + b 60 = -4(5) + b solve b

OpenStudy (kewlgeek555):

\[60=-4(5)+b\]\[60=-20+b\]\[60-(-20)=b\]\[80=b\]

ganeshie8 (ganeshie8):

Yes ! so the equation wud be ?

OpenStudy (kewlgeek555):

c. f(x) = - 4x + 80

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\)

OpenStudy (kewlgeek555):

Thank you so much.

ganeshie8 (ganeshie8):

np :)

OpenStudy (kewlgeek555):

After getting 50/100 on the test two times, I got a 100%!

ganeshie8 (ganeshie8):

you deserve 100% congrats =))

OpenStudy (kewlgeek555):

I meant 90% Im sorry, but still a good grade. ;D

ganeshie8 (ganeshie8):

next time aim for 100% lol